### v1

# An Approximation Algorithm for Monotone Submodular Cost Allocation

Abstract
In this paper, we consider the minimum submodular cost allocation (MSCA) problem. The input of MSCA is $k$ non-negative submodular functions $f_{1},\ldots,f_{k}$ on the ground set $N$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $X_{1},\ldots,X_{k}$ so that $\sum_{i=1}^{k}f_{i}(X_{i})$ is minimized. In this paper, we focus on the case when $f_{1},\ldots,f_{k}$ are monotone (denoted by Mono-MSCA). We provide a natural LP-relaxation for Mono-MSCA, which is equivalent to the convex program relaxation introduced by Chekuri and Ene. We show that the integrality gap of the LP-relaxation is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA. We also show that the integrality gap of the LP-relaxation is at least $k/2-\epsilon$ for any constant $\epsilon>0$ when $k$ is fixed.
1 Introduction
Let $N$ be a finite set of cardinality $n$, and $k\geq 2$ a positive integer. A set function $f:2^{N}\to\mathbb{R}$ is called submodular if $f(S)+f(T)\geq f(S\cup T)+f(S\cap T)$ holds for every $S,T\subseteq N$. In this paper, we focus on the Minimum Submodular Cost Allocation (MSCA) problem, introduced by Chekuri and Ene [5]. The input of MSCA is $k$ non-negative submodular functions $f_{1},\ldots,f_{k}:2^{N}\to\mathbb{R}_{+}$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $X_{1},\ldots,X_{k}$ so that $\sum_{i=1}^{k}f_{i}(X_{i})$ is minimized, i.e.,
$\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
Throughout this paper, we assume that the input functions $f_{1},\ldots,f_{k}$ satisfy $f_{i}(\emptyset)=0$ for each $i=1,\ldots,k$. Several well-studied problems such as the multiway cut problems in graphs and hypergraphs, the submodular multiway partition problems, and uniform metric labeling can be cast as special cases of MSCA. In the case when all submodular functions are monotone, MSCA is equivalent to the submodular facility location problem considered by Svitkina and Tardos [18]. In this paper, we focus on this setting, i.e., the case where each of $f_{1},\ldots,f_{k}$ is monotone submodular. We denote MSCA with monotone submodular functions $f_{1},\ldots,f_{k}$ by Mono-MSCA. The aim of this paper is to investigate the integrality gap of a natural LP-relaxation for Mono-MSCA when $k$ is small.
LP-relaxation.
To provide an LP-relaxation for Mono-MSCA, we introduce some notation. Let $\mathbf{1}\in\mathbb{Z}^{N}_{>0}$ denote the all-ones vector and $k]:=\{1,2,\ldots,k\}$. For any $S\subseteq N$, let $\chi_{S}\in\{0,1\}^{N}$ denote the characteristic vector of $S$. We consider the following linear programming relaxation for MSCA:
$(\textsc{LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
$\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
In this LP-relaxation, we interpret $y_{i}(S)=1$ to mean that the set $S$ is assigned to $f_{i}$ in MSCA, whereas $y_{i}(S)=0$ means that $S$ is not assigned to $f_{i}$. We note that if an integer feasible solution $(y_{1},\ldots,y_{k})$ satisfies $y_{i}(S_{1})=y_{i}(S_{2})=1$ for two distinct sets $S_{1},S_{2}\subseteq N$ and $i\in[k]$, then by the submodularity of $f_{i}$ we can update $y_{i}(S_{1}):=y_{i}(S_{2}):=0$ and $y_{i}(S_{1}\cup S_{2}):=1$ without increasing the objective value. Hence, after such uncrossing operations, any integer feasible solution to LP-Rel induces a partition of $N$ for MSCA.
The linear program LP-Rel is equivalent to the convex program relaxation introduced by Chekuri and Ene [5] (see Appendix A for details). The linear program LP-Rel can also be seen as the dual linear program of the unweighted $k$-polymatroid intersection (see [16] and Appendix B for details). Regarding the integrality gap of LP-Rel, Ene and Vondrák [8] showed that if $\mathrm{P}\neq\mathrm{NP}$, then the integrality gap of LP-Rel for general submodular functions cannot be upper-bounded by any factor. Hence, it is natural to focus on special cases such as a monotone case that admit a bounded integrality gap. We refer to LP-Rel restricted to monotone submodular functions $f_{1},\ldots,f_{k}$ as Mono-LP-Rel. Svitkina and Tardos [18] gave a $(1+\ln n)$-approximation algorithm for Mono-MSCA, and Chekuri and Ene [5] showed that the integrality gap of Mono-LP-Rel is at most $(1+\ln n)$. Santiago and Shepherd [15] gave a $k$-approximation algorithm for Mono-MSCA and showed that the integrality gap of Mono-LP-Rel is at most $k$. On the negative side, Svitkina and Tardos [18] showed that the set cover problem with a collection of $k$ sets can be reduced to Mono-MSCA. This implies that if $k$ is a part of input, Mono-MSCA cannot be approximated within a factor of $(1-o(1))\cdot\ln n$ [10], which matches the approximation ratio shown by Svitkina and Tardos [18]. On the other hand, if $k\leq\log n$, then the set cover problem with a collection of $k$ sets can be solved in time polynomial in $n$ by brute-force. Hence, in this case the reduction does not yield any approximation lower bound for Mono-MSCA. Therefore, the approximability of Mono-MSCA and the integrality gap of Mono-LP-Rel for small $k$ are much less understood. In this paper, we aim to obtain tight upper and lower bounds of the integrality gap of Mono-LP-Rel when $k$ is small.
Our results and techniques.
As mentioned previously, the current best upper bound on the integrality gap of Mono-LP-Rel is $\min\{k,1+\ln n\}$. Our first result improves this to $k/2$.
Theorem 1.1.
The integrality gap of Mono-LP-Rel is at most $k/2$.
When $k=2$, Mono-LP-Rel coincides with the dual linear program of the unweighted polymatroid intersection. Hence, when $k=2$, Theorem 1.1 corresponds to the (total) dual integrality of the polymatroid intersection. Since the proof of Theorem 1.1 is constructive, it yields a $k/2$-approximation algorithm for Mono-MSCA, which improves the current best approximation ratio $\min\{k,1+\ln n\}$.
Theorem 1.2.
There exists a $k/2$-approximation algorithm for Mono-MSCA.
Our next result is a lower bound on the integrality gap of Mono-LP-Rel, which is tight when $k$ is fixed.
Theorem 1.3.
If $\log n\geq k$, then the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
We note that if $k$ is fixed, Theorem 1.3 implies that the integrality gap of Mono-LP-Rel is at least $k/2-\epsilon$ for any constant $\epsilon>0$, which matches the upper bound in Theorem 1.1. Another consequence is that if $k\leq\sqrt{\log n}$, then the lower bound in Theorem 1.3 satisfies $\frac{k}{2}-\frac{k(k-1)}{\log n}\geq\frac{k}{2}-1$, which almost matches the upper bound $k/2$ in Theorem 1.1.
Our proof of Theorem 1.1 is constructive as follows. We first obtain an optimal solution $(y_{1},\ldots,y_{k})$ to Mono-LP-Rel whose supports are chains via the ellipsoid method. Then, we take the top $2/k$ of a fraction of the largest sets in the chains. We show that such a collection of sets can be partitioned into groups of $k$ sets, each of which has union $N$. The existence of this partition implies that the integrality gap of Mono-LP-Rel is at most $k/2$. In the proof of Theorem 1.3, we construct $k$ monotone submodular functions such that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
Related work.
Recently, Bi et al. [3] considered the monotone submodular multiway partition problem (Mono-Sub-MP), which is a special case of Mono-MSCA. They showed that Mono-Sub-MP admits a $4/3$-approximation, and does not admit a $(10/9-\epsilon)$-approximation for every constant $\epsilon>0$. For further recent works related to submodular partition problems, see [1, 2, 4, 11, 14].
There are several recent works related to MSCA. Etesami [9] showed that multi-item welfare maximization under network externalities can be formulated as a special case of MSCA. Deng et al. [7] introduced the generalized load-balancing (GLB) problem, and showed that GLB is equivalent to MSCA when the outer norm is the $\ell_{1}$-norm.
Organization.
The rest of the paper is organized as follows. In Section 2, we show that the integrality gap of Mono-LP-Rel is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA. In Section 3, we show that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$.
2 Approximation algorithm
In this section, we show that the integrality gap of Mono-LP-Rel is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA.
Let $N$ be a finite set of cardinality $n$, and $k\geq 2$ a positive integer. For each $i\in[k]$, let $f_{i}:2^{N}\to\mathbb{R}_{+}$ be a monotone submodular function with $f_{i}(\emptyset)=0$. Recall that Mono-LP-Rel is defined as follows:
$(\textsc{Mono-LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
$\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
To prove the upper bound of the integrality gap, we show that an integer solution to Mono-LP-Rel achieving a $k/2$-approximation can be constructed from an optimal solution to Mono-LP-Rel. Since the coefficient matrix of the constraints in Mono-LP-Rel is rational, Mono-LP-Rel has a rational optimal solution $(y^{*}_{1},\ldots,y^{*}_{k})$. Let $\mathcal{L}_{i}:=\{S\subseteq N\mid y^{*}_{i}(S)>0\}$ and $d_{i}:=|\mathcal{L}_{i}|$ for each $i\in[k]$. By the uncrossing operations, we may assume that $\mathcal{L}_{i}$ is a chain for each $i\in[k]$. Let $C_{i}^{1}\supsetneq C_{i}^{2}\supsetneq\cdots\supsetneq C_{i}^{d_{i}}$ be the members of $\mathcal{L}_{i}$ for each $i\in[k]$. Let $M$ be the product of k(k1)k(k-1) and the denominators of $y^{*}_{i}(C_{i}^{j})$ for all $i,j$. Then, $M\cdot y^{*}_{i}$ is an integer vector for every $i\in[k]$. For each $i\in[k]$ and $j\in[M]$, define $U_{i}^{j}\subseteq N$ as follows:
$U_{i}^{j}:=\begin{cases}C_{i}^{\ell}&\mathrm{if\ }1+\displaystyle\sum_{p=1}^{\ell-1}M\cdot y_{i}^{*}(C_{i}^{p})\leq j\leq\displaystyle\sum_{p=1}^{\ell}M\cdot y_{i}^{*}(C_{i}^{p})\ \mathrm{for\ some\ }\ell\in[d_{i}],\\ \emptyset&\mathrm{if\ }1+\displaystyle\sum_{p=1}^{d_{i}}M\cdot y_{i}^{*}(C_{i}^{p})\leq j\leq M.\end{cases}$
Intuitively, for each $i\in[k]$, we consider the multiset that contains $M\cdot y^{*}_{i}(C_{i}^{\ell})$ copies of $C_{i}^{\ell}$ for all $\ell\in[d_{i}]$, and $U_{i}^{j}$ is the $j$-th largest set in the multiset for each $j\in[M]$. By the definition of $U_{i}^{j}$, we have the following lemma.
Lemma 2.1.
For each $i\in[k]$ and $\ell=1,\ldots,d_{i}$, the number of $j\in[M]$ with $U_{i}^{j}=C_{i}^{\ell}$ is $M\cdot y_{i}^{*}(C_{i}^{\ell})$.
We now aim to construct an integer solution to Mono-LP-Rel (i.e., a partition of $N$) from $U_{i}^{j}$s that achieves $k/2$-approximation. To this end, we prepare the following two lemmas.
Lemma 2.2.
We have
$\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j})=\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S).$
Proof.
By Lemma 2.1, we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{S\in\mathcal{L}_{i},i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{1\leq\ell\leq d_{i},i\in[k]}y^{*}_{i}(C_{i}^{\ell})f_{i}(C_{i}^{\ell})$
$=\sum_{1\leq\ell\leq d_{i},i\in[k]}\dfrac{1}{M}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|\cdot f_{i}(C_{i}^{\ell})$
$=\sum_{\begin{subarray}{c}i\in[k],j\in[M]\\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }1\leq\ell\leq d_{i}\end{subarray}}\dfrac{1}{M}f_{i}(U_{i}^{j})=\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j}).$
The last equality follows since if $U_{i}^{j}\neq C_{i}^{\ell}$ for all $\ell=1,\ldots,d_{i}$, then we have $U_{i}^{j}=\emptyset$ by the definition. ∎
Lemma 2.3.
Let $a_{1},\ldots,a_{k}\in[M]$ be positive integers such that $\sum_{i=1}^{k}a_{i}\leq M+k-1$. Then, $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$.
Proof.
Suppose to the contrary that $\bigcup_{i=1}^{k}U_{i}^{a_{i}}\neq N$. Then, there exists $e\in N$ such that $e\notin U_{i}^{a_{i}}$ for every $i\in[k]$. Combined with $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ for each $i\in[k]$, this implies $e\notin U_{i}^{j}$ for every $i\in[k]$ and integer $j$ with $a_{i}\leq j\leq M$. Hence, by Lemma 2.1 we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(S)=\sum_{\begin{subarray}{c}S\in\mathcal{L}_{i},e\in S\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(S)=\sum_{\begin{subarray}{c}1\leq\ell\leq d_{i},e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(C_{i}^{\ell})$
$=\sum_{\begin{subarray}{c}1\leq\ell\leq d_{i},e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|$
$=\sum_{i\in[k]}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}],\ e\in U_{i}^{j}\right\}\right|$
$\leq\sum_{i\in[k]}\left|\left\{j\in[a_{i}-1]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\right\}\right|$
$\leq\sum_{i\in[k]}(a_{i}-1)\leq M-1.$ (2.1)
Since $y^{*}_{1},\ldots,y^{*}_{k}$ is a feasible solution to Mono-LP-Rel, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y_{i}^{*}(S)=1.$
This implies
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}M\cdot y_{i}^{*}(S)=M,$
which contradicts the inequality (2). ∎
Since $M$ is divisible by k(k1)k(k-1), we can denote $M=k(k-1)\cdot m$ for some integer $m$. For $k$ positive integers $a_{1},\ldots,a_{k}\in[M]$, we say that an ordered tuple $(a_{1},\ldots,a_{k})$ covers $N$ if $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$. We show that if there exist distinct $M\cdot 2/k=2m(k-1)$ ordered tuples covering $N$, then one of them corresponds to an integer solution to Mono-LP-Rel that achieves $k/2$-approximation.
Lemma 2.4.
Suppose that there exist 2m(k1)2m(k-1) ordered tuples
$(a_{1}^{1},\ldots,a_{k}^{1}),(a_{1}^{2},\ldots,a_{k}^{2}),\ldots,(a_{1}^{2m(k-1)},\ldots,a_{k}^{2m(k-1)})$
such that
  • (1)
    $a_{i}^{j}\in[M]$ for each $i\in[k]$ and $j\in[2m(k-1)]$,
  • (2)
    $a_{i}^{j}\neq a_{i}^{\ell}$ for each $i\in[k]$ and $j,\ell\in[2m(k-1)]$ with $j\neq\ell$,
  • (3)
    each ordered tuple covers $N$.
Then, the integrality gap of Mono-LP-Rel is at most $k/2$.
Proof.
By Lemma 2.2, we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j})\geq\sum_{i\in[k],j\in[2m(k-1)]}\dfrac{1}{M}f_{i}(U_{i}^{a_{i}^{j}})$
$\geq\dfrac{2m(k-1)}{M}\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}$
$=\dfrac{2}{k}\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}.$ (2.2)
The first inequality follows since by the monotonicity of $f_{i}$, we have $f_{i}(U_{i}^{j})\geq f_{i}(\emptyset)=0$ for every $i\in[k]$ and $j\in[M]$. Let
$j^{*}:=\underset{j\in[2m(k-1)]}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}).$
Then, by (2) we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\dfrac{2}{k}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$
Hence, we have
$\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$ (2.3)
Since $\bigcup_{i=1}^{k}U_{i}^{a_{i}^{j^{*}}}=N$, there exists a partition $U_{1},\ldots,U_{k}$ of $N$ such that $U_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$. Then, by (2.3) and the monotonicity of $f_{i}$, we have
$\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}})\geq\sum_{i\in[k]}f_{i}(U_{i}).$
This implies that the integrality gap of Mono-LP-Rel is at most $k/2$. ∎
We are now ready to prove Theorem 1.1.
Theorem 1.1.
By Lemma 2.4, it suffices to show that there exist 2m(k1)2m(k-1) ordered tuples satisfying the conditions in Lemma 2.4. For each positive integer $a\in\mathbb{Z}_{>0}$, define $\mathrm{mod}(a)$ as follows:
[…]
For each $i\in[k]$ and $j\in[2m(k-1)]$, define $a_{i}^{j}$ as follows:
$a_{i}^{j}:=\begin{cases}\mathrm{mod}(2m(i-1)+j)&\mathrm{if\ }1\leq i\leq k-1,\\ \mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor&\mathrm{if\ }i=k.\end{cases}$
We show that $(a_{i}^{j})_{i\in[k],j\in[2m(k-1)]}$ satisfies conditions (1)–(3) in Lemma 2.4. For condition (1), we show that $1\leq a_{i}^{j}\leq M$ for each $i\in[k]$ and $j\in[2m(k-1)]$. If $1\leq i\leq k-1$, then we have
$1\leq a_{i}^{j}=\mathrm{mod}(2m(i-1)+j)\leq 2m(k-1)\leq km(k-1)=M.$
If $i=k$, then we have
$a_{i}^{j}=\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor.$
Since (2mj+1)(k1)(2m-j+1)(k-1) is a multiple of k1k-1, we have
$\mathrm{mod}((2m-j+1)(k-1))=p(k-1)$
for some positive integer $p$. This implies
$\mathrm{mod}((2m-j+1)(k-1))\geq k-1.$
Hence, we have
$1=k-1-(k-2)=k-1-\left\lfloor\dfrac{2m(k-1)-1}{2m}\right\rfloor\leq\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor,$
and
$\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor\leq 2m(k-1)\leq km(k-1)=M.$
Therefore, condition (1) holds for $(a_{i}^{j})_{i\in[k],j\in[2m(k-1)]}$.
For condition (2), we next show that $a_{i}^{j}\neq a_{i}^{\ell}$ for each $i\in[k]$ and $j,\ell\in[2m(k-1)]$ with $j\neq\ell$. If $1\leq i\leq k-1$, then by the definition of $a_{i}^{j}$, it is clear that $a_{i}^{j}\neq a_{i}^{\ell}$ holds for each $j,\ell\in[2m(k-1)]$ with $j\neq\ell$. Consider the case when $i=k$. Suppose to the contrary that $a_{k}^{j}=a_{k}^{\ell}$ holds for some $j,\ell\in[2m(k-1)]$ with $j\neq\ell$. This implies that
$\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor=\mathrm{mod}((2m-\ell+1)(k-1))-\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor.$ (2.4)
Since $\mathrm{mod}((2m-j+1)(k-1))$ and $\mathrm{mod}((2m-\ell+1)(k-1))$ are multiples of k1k-1, it follows that
$\left\lfloor\dfrac{j-1}{2m}\right\rfloor-\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor$
is a multiple of k1k-1 by (2.4). Combined with
$0\leq\left\lfloor\dfrac{j-1}{2m}\right\rfloor,\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor\leq k-2,$
this implies that
$\left\lfloor\dfrac{j-1}{2m}\right\rfloor=\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor.$
Hence, we have $2m(p-1)+1\leq j,\ell\leq 2mp$ for some $p\in[k-1]$. Let $j^{\prime}:=2mp-j$ and $\ell^{\prime}:=2mp$-$\ell$. By the definition of $p$, we have $0\leq j^{\prime},\ell^{\prime}\leq 2m-1$. By (2.4), we have
$\mathrm{mod}((2m-j+1)(k-1))=\mathrm{mod}((2m-\ell+1)(k-1)).$ (2.5)
We also have
$\mathrm{mod}((2m-j+1)(k-1))=\mathrm{mod}((2m-2mp+j^{\prime}+1)(k-1))=(j^{\prime}+1)(k-1),$
$\mathrm{mod}((2m-\ell+1)(k-1))=\mathrm{mod}((2m-2mp+\ell^{\prime}+1)(k-1))=(\ell^{\prime}+1)(k-1).$ (2.6)
Hence, by (2.5) and (2) we have
$(j^{\prime}+1)(k-1)=(\ell^{\prime}+1)(k-1).$
This implies $j^{\prime}=\ell^{\prime}$, which contradicts $j\neq\ell$.
For condition (3), we show that $\bigcup_{i=1}^{k}U_{i}^{a_{i}^{j}}=N$ holds for every $j\in[2m(k-1)]$. By Lemma 2.3, it suffices to show that $\sum_{i=1}^{k}a_{i}^{j}\leq M+k-1$ for every $j\in[2m(k-1)]$. Take any $j\in[2m(k-1)]$. Let $p\in[k-1]$ be the unique positive integer such that $2m(p-1)+1\leq j\leq 2mp$. Let $j^{\prime}:=2mp-j$. Then, we have $0\leq j^{\prime}\leq 2m-1$. By the definition of $a_{i}^{j}$, we have
$\sum_{i=1}^{k}a_{i}^{j}=\sum_{i=1}^{k-1}\mathrm{mod}(2m(i-1)+j)+\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor$
$=\sum_{i=1}^{k-1}\mathrm{mod}(2m(i-1)+2mp-j^{\prime})+\mathrm{mod}((2m-2mp+j^{\prime}+1)(k-1))-\left\lfloor\dfrac{2mp-j^{\prime}-1}{2m}\right\rfloor$
$=\sum_{i=1}^{k-p}(2m(i-1)+2mp-j^{\prime})+\sum_{i=k-p+1}^{k-1}(2m(i-1)+2mp-j^{\prime}-2m(k-1))$
$+(j^{\prime}+1)(k-1)-(p-1)$
$=\sum_{i=1}^{k-1}(2m(i-1)+2mp-j^{\prime})-2m(k-1)(p-1)+(j^{\prime}+1)(k-1)-(p-1)$
$=mk(k-1)-(k-1)(2m-2mp+j^{\prime})-2m(k-1)(p-1)+(j^{\prime}+1)(k-1)-(p-1)$
$=M+k-p\leq M+k-1.$
This completes the proof. ∎
The proof of Theorem 1.1 provides the construction of a $k/2$-approximate integer solution to Mono-LP-Rel. We prove Theorem 1.2 by showing that such an integer solution can be constructed in polynomial time.
Theorem 1.2.
We can obtain a rational optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to Mono-LP-Rel whose supports $\mathcal{L}_{1},\ldots,\mathcal{L}_{k}$ are chains in polynomial time via the ellipsoid method [5] (see Appendix A for details). Then, we define $U_{i}^{j}$ for each $i\in[k]$ and $j\in[M]$, and $a_{i}^{j}$ for each $i\in[k]$ and $j\in[2m(k-1)]$ as in the proof of Theorem 1.1. If we can compute
$j^{*}:=\underset{j\in[2m(k-1)]}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
then any partition $U_{1},\ldots,U_{k}$ of $N$ such that $U_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$ satisfies
$\sum_{i\in[k]}f_{i}(U_{i})\leq\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\dfrac{k}{2}\cdot\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\},$
as desired. Hence, it suffices to show that $j^{*}$ can be computed in polynomial time. To find $j^{*}$, we set $j_{1}:=1$. Then, for q=2,3,,q=2,3,..., we define $j_{q}$ as the minimum integer such that $j_{q}>j_{q-1}$ and there exists $i\in[k]$ with $U_{i}^{a_{i}^{j_{q}}}\neq U_{i}^{a_{i}^{j_{q-1}}}$. Recall that $n:=|N|$. Since for each $i\in[k]$, we have $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ and there are at most $|\mathcal{L}_{i}|\leq n$ distinct sets among $\{U_{i}^{j}\}_{j=1}^{M}$, we only need to define $j_{q}$ for $q=1,2,\ldots,Q$ with $Q\leq kn$. Since we have
$j^{*}=\underset{j\in\{j_{1},j_{2},\ldots,j_{Q}\}}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
we can find $j^{*}$ by computing the values of $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$. By the definition of $U_{i}^{j}$ and $a_{i}^{j}$, we can compute $j_{q}$ in polynomial time for each fixed $q$. Hence, since $Q\leq kn$, we can compute $j_{q}$ and $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$ in polynomial time. ∎
3 Lower bound on the integrality gap
In this section, we show that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$. To this end, we explicitly construct monotone submodular functions $f_{1},\ldots,f_{k}$ such that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
Suppose that $\log n\geq k$. Let $p$ be the unique positive integer such that $pk\leq\lfloor\log n\rfloor\leq(p+1)k-1$. We define
$\mathcal{V}:=\left\{v\in\mathbb{Z}_{+}^{k}\ \middle|\ \sum_{i=1}^{k}v_{i}=pk-k+1\right\}.$
By stars and bars (see e.g., [17, Chapter 1]), we have
$|\mathcal{V}|=\binom{pk}{k-1}\leq 2^{pk}\leq 2^{\lfloor\log n\rfloor}\leq n=|N|.$
Choose $e_{v}\in N$ for each $v\in\mathcal{V}$ so that all $e_{v}$ are distinct. For each $i\in[k]$, let
$N_{i}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}\geq 1\},$
$Z_{i}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}=0\}.$
For each $i\in[k]$, define $f_{i}:2^{N}\to\mathbb{R}$ as follows: for all $S\subseteq N$,
$f_{i}(S):=\begin{cases}\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}\}&\mathrm{if\ }S\cap N_{i}\neq\emptyset\ \mathrm{and\ }S\cap Z_{i}=\emptyset,\\ 0&\mathrm{if\ }S\cap N_{i}=\emptyset\ \mathrm{and\ }S\cap Z_{i}=\emptyset,\\ 2pk+1&\mathrm{if\ }S\cap Z_{i}\neq\emptyset.\end{cases}$
Lemma 3.1.
$f_{i}$ is monotone submodular for each $i\in[k]$.
Proof.
We first show monotonicity of $f_{i}$. Take any $S,T\subseteq N$ with $S\subseteq T$. If $S\cap N_{i}=S\cap Z_{i}=\emptyset$, then we have $f_{i}(S)=0\leq f_{i}(T)$. If $S\cap Z_{i}\neq\emptyset$, then we have $T\cap Z_{i}\neq\emptyset$, which implies $f_{i}(S)=f_{i}(T)=2pk+1$. Consider the case when $S\cap N_{i}\neq\emptyset$ and $S\cap Z_{i}=\emptyset$. If $T\cap Z_{i}\neq\emptyset$, then we have
$f_{i}(S)=\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}\}\leq 2p+1-1=2p\leq 2pk+1=f_{i}(T).$
If $T\cap Z_{i}=\emptyset$, then since $T\cap N_{i}\neq\emptyset$ follows from $S\cap N_{i}\neq\emptyset$, we have
$f_{i}(S)$ $=\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}\}$
$\leq\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in T\cap N_{i}\}\}=f_{i}(T).$
Therefore, $f_{i}$ is monotone.
We next show that $f_{i}$ is submodular. Take any $S,T\subseteq N$. If $S\cap Z_{i}\neq\emptyset$, then by the monotonicity of $f_{i}$, we have
$f_{i}(S)+f_{i}(T)=2pk+1+f_{i}(T)\geq 2pk+1+f_{i}(S\cap T)\geq f_{i}(S\cup T)+f_{i}(S\cap T).$
The last inequality follows since $f_{i}(X)\leq 2pk+1$ holds for every $X\subseteq N$ by the definition of $f_{i}$. If $S\cap N_{i}=S\cap Z_{i}=\emptyset$, then since $(S\cup T)\cap N_{i}=T\cap N_{i}$ and $(S\cup T)\cap Z_{i}=T\cap Z_{i}$, we have $f_{i}(S\cup T)=f_{i}(T)$ by the definition of $f_{i}$. This implies that
$f_{i}(S)+f_{i}(T)=f_{i}(S)+f_{i}(S\cup T)\geq f_{i}(S\cup T)+f_{i}(S\cap T).$
The last inequality follows from the monotonicity of $f_{i}$. Hence, by the symmetry between $S$ and $T$, the remaining case is when $S\cap N_{i}\neq\emptyset$ and $S\cap Z_{i}=\emptyset$, and $T\cap N_{i}\neq\emptyset$ and $T\cap Z_{i}=\emptyset$. In this case, we have $(S\cup T)\cap N_{i}\neq\emptyset$ and $(S\cup T)\cap Z_{i}=\emptyset$, which implies that
$f_{i}(S\cup T)=\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in(S\cup T)\cap N_{i}\}\}.$
Take $e_{v^{*}}\in\mathrm{argmax}\{2p+1-v_{i}\mid e_{v}\in(S\cup T)\cap N_{i}\}$ with the corresponding vector $v^{*}\in\mathcal{V}$. Then, we have $e_{v^{*}}\in S\cap N_{i}$ or $e_{v^{*}}\in T\cap N_{i}$. We may assume without loss of generality that $e_{v^{*}}\in S\cap N_{i}$. Then, we have $e_{v^{*}}\in\mathrm{argmax}\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}$. Hence, we have
$f_{i}(S)+f_{i}(T)=\max\{0,2p+1-v^{*}_{i}\}+f_{i}(T)=f_{i}(S\cup T)+f_{i}(T)$ $\geq f_{i}(S\cup T)+f_{i}(S\cap T).$
The last inequality follows from the monotonicity of $f_{i}$. Therefore, $f_{i}$ is submodular. ∎
We now show that for the monotone submodular functions $f_{1},\ldots,f_{k}$ defined above, any optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to Mono-LP-Rel satisfies
$\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$
$\leq\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$ (3.1)
The inequality (3) implies that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$, which completes the proof of Theorem 1.3.
To prove (3), we first give an upper bound on the value $\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$. To this end, we construct a feasible solution to Mono-LP-Rel. For each $i\in[k]$ and $j\in[pk-k+1]$, let $N_{i}^{j}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}\geq j\}$. Note that $N_{i}^{1}=N_{i}$. In what follows, we will use $N_{i}^{1}$ instead of $N_{i}$ for convenience. Let $N^{c}:=N\setminus\{e_{v}\in N\mid v\in\mathcal{V}\}$. For each $i\in[k]$, define $z_{i}\in\mathbb{R}_{+}^{2^{N}}$ as follows: for all $S\subseteq N$,
$z_{i}(S):=\begin{cases}\dfrac{1}{pk-k+1}&\mathrm{if\ }i=1\ \mathrm{and\ }\vskip 3.0ptS=N_{1}^{j}\cup N^{c}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ \vskip 3.0pt\dfrac{1}{pk-k+1}&\mathrm{if\ }2\leq i\leq k\ \mathrm{and\ }S=N_{i}^{j}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ 0&\mathrm{otherwise}.\end{cases}$
Lemma 3.2.
$(z_{1},\ldots,z_{k})$ is a feasible solution to Mono-LP-Rel.
Proof.
To show the feasibility of $(z_{1},\ldots,z_{k})$, we show that
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=1$ (3.2)
for every $e\in N$. For any $e\in N^{c}$, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=\sum_{j\in[pk-k+1]}z_{1}(N_{1}^{j}\cup N^{c})=\sum_{j\in[pk-k+1]}\dfrac{1}{pk-k+1}=1.$
Take any $e\in N\setminus N^{c}$. Since $e=e_{v}$ for some $v\in\mathcal{V}$, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)$ $=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}z_{1}(N_{1}^{j}\cup N^{c})+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}z_{i}(N_{i}^{j})$
$=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}\frac{1}{pk-k+1}+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}\frac{1}{pk-k+1}$
$=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ i\in[k],v_{i}\geq j\end{subarray}}\frac{1}{pk-k+1}$
$=\frac{pk-k+1}{pk-k+1}=1.$
The second last equality follows since $\sum_{i=1}^{k}v_{i}=pk-k+1$. Hence, (3.2) holds for every $e\in N$, which implies that $(z_{1},\ldots,z_{k})$ is a feasible solution to Mono-LP-Rel. ∎
By Lemma 3.2 and the optimality of $y^{*}_{1},\ldots,y^{*}_{k}$, we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)$
$=\sum_{j\in[pk-k+1]}\frac{1}{pk-k+1}f_{1}(N_{1}^{j}\cup N^{c})+\sum_{2\leq i\leq k,j\in[pk-k+1]}\frac{1}{pk-k+1}f_{i}(N_{i}^{j})$
$=\sum_{j\in[pk-k+1]}\frac{1}{pk-k+1}\cdot\max\{0,2p+1-j\}+\sum_{2\leq i\leq k,j\in[pk-k+1]}\frac{1}{pk-k+1}\cdot\max\{0,2p+1-j\}$
$=\frac{pk(2p+1)}{pk-k+1}.$ (3.3)
To prove (3), we next provide a lower bound on the right hand side of (3).
Lemma 3.3.
Let $X_{1},\ldots,X_{k}$ be any partition of $N$. Then, we have
$\sum_{i=1}^{k}f_{i}(X_{i})\geq pk+k.$
Proof.
If $X_{j}\cap Z_{j}\neq\emptyset$ for some $j\in[k]$, then we have
$\sum_{i=1}^{k}f_{i}(X_{i})\geq f_{j}(X_{j})=2pk+1\geq pk+k.$
Hence, it suffices to consider the case when $X_{i}\cap Z_{i}=\emptyset$ for every $i\in[k]$. For each $i\in[k]$, define $x^{i}\in\mathcal{V}$ as follows:
$x^{i}_{j}:=\begin{cases}pk-k+1&\mathrm{if\ }j=i,\\ 0&\mathrm{if\ }j\in[k]\mathrm{\ and\ }j\neq i.\end{cases}$
Then, since $X_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have $e_{x^{i}}\in X_{i}$ for every $i\in[k]$. Hence, we have $X_{i}\cap N_{i}\neq\emptyset$ for every $i\in[k]$. For each $i\in[k]$, we define $b_{i}:=\min\{v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}$. We now show that $\sum_{i=1}^{k}b_{i}\leq pk$. Suppose to the contrary that $\sum_{i=1}^{k}b_{i}\geq pk+1$. Then, since $\sum_{i=1}^{k}(b_{i}-1)\geq pk-k+1$, there exists $v\in\mathcal{V}$ such that $v_{i}\leq b_{i}-1$ for every $i\in[k]$. Choose $i\in[k]$ such that $e_{v}\in X_{i}$. Since $X_{i}\cap Z_{i}=\emptyset$ by the assumption, we have $e_{v}\in X_{i}\cap N_{i}$, which contradicts the definition of $b_{i}$. Hence, we have $\sum_{i=1}^{k}b_{i}\leq pk$. Therefore, since $X_{i}\cap N_{i}\neq\emptyset$ and $X_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have
$\sum_{i=1}^{k}f_{i}(X_{i})=\sum_{i=1}^{k}\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}\}$
$\geq\sum_{i=1}^{k}\max\{2p+1-v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}=2pk+k-\sum_{i=1}^{k}\min\{v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}$
$=2pk+k-\sum_{i=1}^{k}b_{i}\geq 2pk+k-pk=pk+k.$
By Lemma 3.3, we have
$\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}$
$\geq pk+k=\frac{2k(p+1)^{2}(pk-k+1)}{2(p+1)(pk-k+1)}$
$=\frac{2k(p+1)(p^{2}k-k+p+1)}{2(p+1)(pk-k+1)}=\frac{2k(p+1)(p^{2}k+(p-1)(k-1)-pk+2p)}{2(p+1)(pk-k+1)}$
$\geq\frac{2k(p+1)(p^{2}k-pk+2p)}{2(p+1)(pk-k+1)}=\frac{pk(2p+2)(pk-k+2)}{2(p+1)(pk-k+1)}$
$\geq\frac{pk(2p+1)(pk-k+2)}{2(p+1)(pk-k+1)}=\frac{pk(2p+1)}{pk-k+1}\cdot\left(\frac{k}{2}-\frac{k(k-1)}{k(p+1)}\right)$
$\geq\frac{pk(2p+1)}{pk-k+1}\cdot\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right).$ (3.4)
The last inequality follows from the definition of $p$. Combining (3) and (3), we have
$\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\frac{pk(2p+1)}{pk-k+1}$
$\leq\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
Thus, the inequality (3) holds, which completes the proof of Theorem 1.3.
Acknowledgements
This work was supported by JST ERATO Grant Number JPMJER2301, Japan.
Appendix A Convex program based on the Lovász extension
In this section, we show that LP-Rel is equivalent to the convex program relaxation of MSCA introduced by Chekuri and Ene [5]. Let $N:=\{1,\ldots,n\}$. Let $f_{i}:2^{N}\to\mathbb{R}$ be a submodular function with $f_{i}(\emptyset)=0$ for each $i\in[k]$. Recall that LP-Rel is defined as follows:
$(\textsc{LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
$\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
For a set function $f:2^{N}\to\mathbb{R}$ with $f(\emptyset)=0$, the Lovász extension $\hat{f}:[0,1]^{n}\to\mathbb{R}$ of $f$ is defined as follows: Take any $x\in[0,1]^{n}$. Relabel the elements in $N$ so that $x_{1}\geq x_{2}\geq\cdots\geq x_{n}$. For convenience, let $x_{n+1}:=0$. Let $S_{i}:=\{1,2,\ldots,i\}$ for each $i\in[n]$. The value of $\hat{f}(x)$ is defined as follows:
$\hat{f}(x):=\sum_{i=1}^{n}(x_{i}-x_{i+1})f(S_{i}).$
Chekuri and Ene [5] introduced the following convex program relaxation of MSCA based on the Lovász extension:
$(\textsc{LE-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{i=1}^{k}\hat{f}_{i}(x^{i})$
$\mathrm{subject\ to}$ $\ \ \ x^{i}\in[0,1]^{n}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{i=1}^{k}x^{i}_{j}=1\ \mathrm{for\ every\ }j\in[n].$
We now show that any feasible solution $({x^{1}},\ldots,x^{k})$ to LE-Rel can be translated into a feasible solution $(y_{1},\ldots,y_{k})$ to LP-Rel that has the same objective value and whose supports are chains. Take any feasible solution $({x^{1}},\ldots,x^{k})$ to LE-Rel. We will describe a translation from $x^{i}$ to $y_{i}$ for each $i\in[k]$. We first relabel the elements in $N$ so that $x^{i}_{1}\geq x^{i}_{2}\geq\cdots\geq x^{i}_{n}$. Let $x^{i}_{n+1}:=0$ and $S_{j}:=\{1,2,\ldots,j\}$ for each $j\in[n]$. Then, for all $S\subseteq N$ define $y_{i}(S)$ as follows:
$y_{i}(S):=\begin{cases}x_{j}^{i}-x_{j+1}^{i}&\mathrm{if\ }S=S_{j}\ \mathrm{and\ }x_{j}^{i}-x_{j+1}^{i}>0\ \mathrm{for\ some\ }j\in[n],\\ 0&\mathrm{otherwise}.\end{cases}$
Then, $(y_{1},\ldots,y_{k})$ is a feasible solution to LP-Rel because for each $i\in[k]$ we have
$\sum_{S\subseteq N}y_{i}(S)\chi_{S}=\sum_{j\in[n]}(x^{i}_{j}-x^{i}_{j+1})\chi_{S_{j}}=x^{i}$
and
$\sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\sum_{i\in[k]}x^{i}=\mathbf{1}.$
For each $i\in[k]$, we also have
$\sum_{S\subseteq N}y_{i}(S)f_{i}(S)=\sum_{j\in[n]}(x_{j}^{i}-x_{j+1}^{i})f_{i}(S_{j})=\hat{f_{i}}(x^{i}).$
Hence, we obtain
$\sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)=\sum_{i=1}^{k}\hat{f_{i}}(x^{i}),$
which implies that $(y_{1},\ldots,y_{k})$ has the same objective value as $(x^{1},\ldots,x^{k})$. Similarly, we can translate any feasible solution to LP-Rel whose supports are chains into a feasible solution to LE-Rel that has the same objective value.
Since LE-Rel can be solved in polynomial time via the ellipsoid method [5], an optimal solution to LP-Rel whose supports are chains can be obtained in polynomial time. Chekuri and Ene [6] provided another equivalent linear programming formulation of MSCA called LE-Rel-Primal. We note that LP-Rel is equivalent to LE-Rel-Primal.
Appendix B $k$-polymatroid intersection
Let $N$ be a finite set and $k\geq 2$ a positive integer. For each $i\in[k]$, let $f_{i}:2^{N}\to\mathbb{R}$ be a submodular function with $f_{i}(\emptyset)=0$. For a set function $f:2^{N}\to\mathbb{R}$, we define $P(f):=\{x\in\mathbb{R}^{N}\mid x(S)\leq f(S)\ \mathrm{for\ every\ }S\subseteq N\}$, where $x(S):=\sum_{e\in S}x(e)$ for every $S\subseteq N$. Let $w\in\mathbb{R}_{+}^{N}$ be a weight vector. The weighted $k$-polymatroid intersection problem is defined as follows:
$(\mathrm{LP}_{k\text{-poly}})\ \ \ \ \ \mathrm{maximize}$ $\ \ \ w^{\top}x$
$\mathrm{subject\ to}$ $\ \ \ x\in P(f_{i})\ \mathrm{for\ every\ }i\in[k].$
The linear program $\mathrm{LP}_{k\text{-poly}}$ corresponds to the weighted polymatroid intersection when $k=2$. When $w=\mathbf{1}$, the linear program $\mathrm{LP}_{k\text{-poly}}$ is the dual linear program of LP-Rel. For the case when $f_{1},\ldots,f_{k}$ are matroid rank functions (i.e., the case of weighted $k$-matroid intersection), the integrality gap of $\mathrm{LP}_{k\text{-poly}}$ is $2$ when $k=3$ [13], and it equals k1k-1 for arbitrary $k$ in the unweighted case $w=\mathbf{1}$ [12].

### v2

@@
 In this paper, we consider the minimum submodular cost allocation (MSCA) problem.
-The input of MSCA is $k$ non-negative submodular functions $f_{1},\ldots,f_{k}$ on the ground set $N$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $X_{1},\ldots,X_{k}$ so that $\sum_{i=1}^{k}f_{i}(X_{i})$ is minimized.
-In this paper, we focus on the case when $f_{1},\ldots,f_{k}$ are monotone (denoted by Mono-MSCA).
-We provide a natural LP-relaxation for Mono-MSCA, which is equivalent to the convex program relaxation introduced by Chekuri and Ene.
-We show that the integrality gap of the LP-relaxation is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA.
-We also show that the integrality gap of the LP-relaxation is at least $k/2-\epsilon$ for any constant $\epsilon>0$ when $k$ is fixed.
+The input of MSCA is $k$ non-negative submodular functions $f_{1},f_{2},\ldots,f_{k}$ on the ground set $N$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $S_{1},S_{2},\ldots,S_{k}$ so that $\sum_{i=1}^{k}f_{i}(S_{i})$ is minimized.
+In this paper, we focus on the case when $f_{1},f_{2},\ldots,f_{k}$ are monotone, which coincides with the submodular facility location problem considered by Svitkina and Tardos.
+We show that the integrality gap of a natural LP-relaxation for MSCA with monotone submodular functions is at most $k/2$, yielding a $k/2$-approximation algorithm.
+We also prove a nearly matching lower bound: the integrality gap is at least $k/2-\epsilon$ for any constant $\epsilon>0$ when $k$ is fixed.
 1 Introduction
-Let $N$ be a finite set of cardinality $n$, and $k\geq 2$ a positive integer.
+Let $N$ be a finite set and $k\geq 2$ a positive integer.
 A set function $f:2^{N}\to\mathbb{R}$ is called submodular if $f(S)+f(T)\geq f(S\cup T)+f(S\cap T)$ holds for every $S,T\subseteq N$.
 In this paper, we focus on the Minimum Submodular Cost Allocation (MSCA) problem, introduced by Chekuri and Ene [5].
-The input of MSCA is $k$ non-negative submodular functions $f_{1},\ldots,f_{k}:2^{N}\to\mathbb{R}_{+}$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $X_{1},\ldots,X_{k}$ so that $\sum_{i=1}^{k}f_{i}(X_{i})$ is minimized, i.e.,
-$\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
-Throughout this paper, we assume that the input functions $f_{1},\ldots,f_{k}$ satisfy $f_{i}(\emptyset)=0$ for each $i=1,\ldots,k$.
-Several well-studied problems such as the multiway cut problems in graphs and hypergraphs, the submodular multiway partition problems, and uniform metric labeling can be cast as special cases of MSCA.
-In the case when all submodular functions are monotone, MSCA is equivalent to the submodular facility location problem considered by Svitkina and Tardos [18].
-In this paper, we focus on this setting, i.e., the case where each of $f_{1},\ldots,f_{k}$ is monotone submodular.
-We denote MSCA with monotone submodular functions $f_{1},\ldots,f_{k}$ by Mono-MSCA.
-The aim of this paper is to investigate the integrality gap of a natural LP-relaxation for Mono-MSCA when $k$ is small.
+The input of MSCA is $k$ non-negative submodular functions $f_{1},f_{2},\ldots,f_{k}:2^{N}\to\mathbb{R}_{+}$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $S_{1},S_{2},\ldots,S_{k}$ so that $\sum_{i=1}^{k}f_{i}(S_{i})$ is minimized, i.e.,
+$\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
+Several well-studied problems such as the multiway cut problems in graphs and hypergraphs, the submodular multiway partition problems, and uniform metric labeling can be cast as special cases of MSCA [5].
+When all submodular functions are monotone, MSCA is equivalent to the submodular facility location problem considered by Svitkina and Tardos [18].
+In this paper, we focus on this setting, that is, we assume that $f_{i}(S)\leq f_{i}(T)$ for all $S\subseteq T\subseteq N$ and $i=1,2,\ldots,k$.
+We call this monotone case of MSCA the Minimum Monotone Submodular Cost Allocation (MMSCA) problem.
+The aim of this paper is to investigate the integrality gap of a natural LP-relaxation for MMSCA when $k$ is small.
 LP-relaxation.
-To provide an LP-relaxation for Mono-MSCA, we introduce some notation.
+To provide an LP-relaxation for MMSCA, we introduce some notation.
 Let $\mathbf{1}\in\mathbb{Z}^{N}_{>0}$ denote the all-ones vector and $k]:=\{1,2,\ldots,k\}$.
 For any $S\subseteq N$, let $\chi_{S}\in\{0,1\}^{N}$ denote the characteristic vector of $S$.
+Throughout this paper, we assume that the input functions $f_{1},f_{2},\ldots,f_{k}$ satisfy $f_{i}(\emptyset)=0$ for each $i=1,2,\ldots,k$.
 We consider the following linear programming relaxation for MSCA:
-$(\textsc{LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
+$(\textsc{LP})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
 $\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
 $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
-In this LP-relaxation, we interpret $y_{i}(S)=1$ to mean that the set $S$ is assigned to $f_{i}$ in MSCA, whereas $y_{i}(S)=0$ means that $S$ is not assigned to $f_{i}$.
-We note that if an integer feasible solution $(y_{1},\ldots,y_{k})$ satisfies $y_{i}(S_{1})=y_{i}(S_{2})=1$ for two distinct sets $S_{1},S_{2}\subseteq N$ and $i\in[k]$, then by the submodularity of $f_{i}$ we can update $y_{i}(S_{1}):=y_{i}(S_{2}):=0$ and $y_{i}(S_{1}\cup S_{2}):=1$ without increasing the objective value.
-Hence, after such uncrossing operations, any integer feasible solution to LP-Rel induces a partition of $N$ for MSCA.
-The linear program LP-Rel is equivalent to the convex program relaxation introduced by Chekuri and Ene [5] (see Appendix A for details).
-The linear program LP-Rel can also be seen as the dual linear program of the unweighted $k$-polymatroid intersection (see [16] and Appendix B for details).
-Regarding the integrality gap of LP-Rel, Ene and Vondrák [8] showed that if $\mathrm{P}\neq\mathrm{NP}$, then the integrality gap of LP-Rel for general submodular functions cannot be upper-bounded by any factor.
+In this LP-relaxation, we interpret $y_{i}(S)=1$ to mean that $S$ is assigned to $f_{i}$ in MSCA, whereas $y_{i}(S)=0$ means that $S$ is not assigned to $f_{i}$.
+We note that if an integer feasible solution $(y_{1},y_{2},\ldots,y_{k})$ satisfies $y_{i}(S)=y_{i}(T)=1$ for two distinct sets $S,T\subseteq N$ and $i\in[k]$, then by the submodularity of $f_{i}$ we can update $y_{i}(S):=y_{i}(T):=0$ and $y_{i}(S\cup T):=1$ without increasing the objective value.
+Hence, after such uncrossing operations, any integer feasible solution to (LP) induces a partition of $N$ for MSCA.
+The linear program (LP) is equivalent to the convex program relaxation introduced by Chekuri and Ene [5].
+The linear program (LP) can also be seen as the dual linear program of the unweighted $k$-polymatroid intersection; see, e.g., [16].
+Regarding the integrality gap of (LP), Ene and Vondrák [8] showed that if $\mathrm{P}\neq\mathrm{NP}$, then the integrality gap of (LP) for general submodular functions cannot be upper-bounded by any factor.
 Hence, it is natural to focus on special cases such as a monotone case that admit a bounded integrality gap.
-We refer to LP-Rel restricted to monotone submodular functions $f_{1},\ldots,f_{k}$ as Mono-LP-Rel.
-Svitkina and Tardos [18] gave a $(1+\ln n)$-approximation algorithm for Mono-MSCA, and Chekuri and Ene [5] showed that the integrality gap of Mono-LP-Rel is at most $(1+\ln n)$.
-Santiago and Shepherd [15] gave a $k$-approximation algorithm for Mono-MSCA and showed that the integrality gap of Mono-LP-Rel is at most $k$.
-On the negative side, Svitkina and Tardos [18] showed that the set cover problem with a collection of $k$ sets can be reduced to Mono-MSCA.
-This implies that if $k$ is a part of input, Mono-MSCA cannot be approximated within a factor of $(1-o(1))\cdot\ln n$ [10], which matches the approximation ratio shown by Svitkina and Tardos [18].
+We refer to (LP) restricted to monotone submodular functions $f_{1},f_{2},\ldots,f_{k}$ as $\mathrm{(LP_{mono})}$.
+Svitkina and Tardos [18] gave a $(1+\ln n)$-approximation algorithm for MMSCA, and Chekuri and Ene [5] showed that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $(1+\ln n)$, where $n:=|N|$.
+Santiago and Shepherd [15] showed that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k$, yielding a $k$-approximation for MMSCA.
+On the negative side, Svitkina and Tardos [18] showed that the set cover problem with a collection of $k$ sets can be reduced to MMSCA.
+This implies that if $k$ is a part of input, MMSCA cannot be approximated within a factor of $(1-o(1))\cdot\ln n$ [10], which matches the approximation ratio shown by Svitkina and Tardos [18].
 On the other hand, if $k\leq\log n$, then the set cover problem with a collection of $k$ sets can be solved in time polynomial in $n$ by brute-force.
-Hence, in this case the reduction does not yield any approximation lower bound for Mono-MSCA.
-Therefore, the approximability of Mono-MSCA and the integrality gap of Mono-LP-Rel for small $k$ are much less understood.
-In this paper, we aim to obtain tight upper and lower bounds of the integrality gap of Mono-LP-Rel when $k$ is small.
+Hence, in this case the reduction does not yield any approximation lower bound for MMSCA.
+Therefore, the approximability of MMSCA and the integrality gap of $\mathrm{(LP_{mono})}$ for small $k$ are much less understood.
+In this paper, we aim to obtain tight upper and lower bounds of the integrality gap of $\mathrm{(LP_{mono})}$ when $k$ is small.
 Our results and techniques.
-As mentioned previously, the current best upper bound on the integrality gap of Mono-LP-Rel is $\min\{k,1+\ln n\}$. Our first result improves this to $k/2$.
+As mentioned previously, the current best upper bound on the integrality gap of $\mathrm{(LP_{mono})}$ is $\min\{k,1+\ln n\}$. Our first result improves this to $k/2$.
 Theorem 1.1.
-The integrality gap of Mono-LP-Rel is at most $k/2$.
-When $k=2$, Mono-LP-Rel coincides with the dual linear program of the unweighted polymatroid intersection.
-Hence, when $k=2$, Theorem 1.1 corresponds to the (total) dual integrality of the polymatroid intersection.
-Since the proof of Theorem 1.1 is constructive, it yields a $k/2$-approximation algorithm for Mono-MSCA, which improves the current best approximation ratio $\min\{k,1+\ln n\}$.
+The integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$.
+Recall that when $k=2$, $\mathrm{(LP_{mono})}$ is the dual linear program of the unweighted polymatroid intersection.
+Since the constraint system of the unweighted polymatroid intersection is totally dual integral (see, e.g., [16]), its dual admits an integral optimal solution, which matches Theorem 1.1 for $k=2$.
+Hence, Theorem 1.1 can be viewed as extending the (total) dual integrality of polymatroid intersection, in the sense that it yields a $k/2$ bound on the dual integrality gap for $k$-polymatroid intersection.
+Since the proof of Theorem 1.1 is constructive, it yields a $k/2$-approximation algorithm for MMSCA, which improves the current best approximation ratio $\min\{k,1+\ln n\}$.
 Theorem 1.2.
-There exists a $k/2$-approximation algorithm for Mono-MSCA.
-Our next result is a lower bound on the integrality gap of Mono-LP-Rel, which is tight when $k$ is fixed.
+There exists a $k/2$-approximation algorithm for MMSCA.
+Our next result is a lower bound on the integrality gap of $\mathrm{(LP_{mono})}$, which is tight when $k$ is fixed.
 Theorem 1.3.
-If $\log n\geq k$, then the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
-We note that if $k$ is fixed, Theorem 1.3 implies that the integrality gap of Mono-LP-Rel is at least $k/2-\epsilon$ for any constant $\epsilon>0$, which matches the upper bound in Theorem 1.1.
+If $\log n\geq k$, then the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
+We note that if $k$ is fixed, Theorem 1.3 implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $k/2-\epsilon$ for any constant $\epsilon>0$, which matches the upper bound in Theorem 1.1.
 Another consequence is that if $k\leq\sqrt{\log n}$, then the lower bound in Theorem 1.3 satisfies $\frac{k}{2}-\frac{k(k-1)}{\log n}\geq\frac{k}{2}-1$, which almost matches the upper bound $k/2$ in Theorem 1.1.
 Our proof of Theorem 1.1 is constructive as follows.
-We first obtain an optimal solution $(y_{1},\ldots,y_{k})$ to Mono-LP-Rel whose supports are chains via the ellipsoid method.
+We first obtain an optimal solution $(y_{1},y_{2},\ldots,y_{k})$ to $\mathrm{(LP_{mono})}$ whose supports are chains via the ellipsoid method.
 Then, we take the top $2/k$ of a fraction of the largest sets in the chains.
 We show that such a collection of sets can be partitioned into groups of $k$ sets, each of which has union $N$.
-The existence of this partition implies that the integrality gap of Mono-LP-Rel is at most $k/2$.
-In the proof of Theorem 1.3, we construct $k$ monotone submodular functions such that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
+The existence of this partition implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$.
+In the proof of Theorem 1.3, we construct $k$ monotone submodular functions such that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
 Related work.
+Svitkina and Tardos [18] considered the submodular facility location problem defined as follows.
+The input consists of a set of facilities $\mathcal{F}$, a set of demands or clients $\mathcal{D}$, a non-negative cost $c_{ij}$ to connect facility $i$ to client $j$ for each $i\in\mathcal{F}$ and $j\in\mathcal{D}$, and a facility opening cost function $f_{i}$ for each $i\in\mathcal{F}$.
+The goal is to assign each client to a facility so as to minimize the sum of the facility opening cost and the connection cost, where the facility opening cost for a facility $i$ is a monotone submodular function $f_{i}$ of the clients assigned to $i$.
+Since the sum of the facility opening cost and the connection cost for a facility $i$ is a monotone submodular function of the clients assigned to $i$, the submodular facility location problem is equivalent to MMSCA.
+As mentioned above, Svitkina and Tardos [18] gave a $(1+\ln|\mathcal{D}|)$-approximation algorithm for the submodular facility location problem, and showed a matching lower bound $(1-o(1))\cdot\ln|\mathcal{D}|$ when $|\mathcal{F}|$ is not fixed or bounded, while our result improves the approximation ratio to $|\mathcal{F}|/2$.
 Recently, Bi et al.
-[3] considered the monotone submodular multiway partition problem (Mono-Sub-MP), which is a special case of Mono-MSCA.
-They showed that Mono-Sub-MP admits a $4/3$-approximation, and does not admit a $(10/9-\epsilon)$-approximation for every constant $\epsilon>0$.
+[3] considered the submodular multiway partition problem for monotone submodular functions, which they call monotone submodular multiway partition (Mono-Sub-MP).
+In Mono-Sub-MP, we are given a non-negative monotone submodular function $f:2^{N}\to\mathbb{R}_{+}$ with $f(\emptyset)=0$ and $k$ terminals $t_{1},t_{2},\ldots,t_{k}\in N$.
+The goal is to find a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ with $t_{i}\in S_{i}$ for every $i\in[k]$ in order to minimize $\sum_{i=1}^{k}f(S_{i})$.
+As the submodular multiway partition problem can be seen as a special case of MSCA, Mono-Sub-MP is a special case of MMSCA as follows.
+Define $N^{\prime}:=N\setminus\{t_{1},t_{2},\ldots,t_{k}\}$ and $f_{i}(S):=f(S\cup\{t_{i}\})-f(\{t_{i}\})$ for each $i\in[k]$ and $S\subseteq N^{\prime}$.
+Then, $f_{i}$ is a non-negative monotone submodular functions with $f_{i}(\emptyset)=0$ for every $i\in[k]$.
+Hence, Mono-Sub-MP can be cast as a special case of MMSCA.
+Bi et al.
+[3] showed that Mono-Sub-MP admits a $4/3$-approximation, and does not admit a $(10/9-\epsilon)$-approximation for every constant $\epsilon>0$.
 For further recent works related to submodular partition problems, see [1, 2, 4, 11, 14].
@@
 The rest of the paper is organized as follows.
-In Section 2, we show that the integrality gap of Mono-LP-Rel is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA.
-In Section 3, we show that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$.
+In Section 2, we show that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$, which yields a $k/2$-approximation algorithm for MMSCA.
+In Section 3, we show that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$.
 2 Approximation algorithm
-In this section, we show that the integrality gap of Mono-LP-Rel is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA.
+In this section, we prove Theorems 1.1 and 1.2, showing that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$, which yields a $k/2$-approximation algorithm for MMSCA.
 Let $N$ be a finite set of cardinality $n$, and $k\geq 2$ a positive integer.
 For each $i\in[k]$, let $f_{i}:2^{N}\to\mathbb{R}_{+}$ be a monotone submodular function with $f_{i}(\emptyset)=0$.
-Recall that Mono-LP-Rel is defined as follows:
-$(\textsc{Mono-LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
+Recall that $\mathrm{(LP_{mono})}$ is defined as follows:
+$\textsc{$\mathrm{(LP_{mono})}$}\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
 $\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
 $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
-To prove the upper bound of the integrality gap, we show that an integer solution to Mono-LP-Rel achieving a $k/2$-approximation can be constructed from an optimal solution to Mono-LP-Rel.
-Since the coefficient matrix of the constraints in Mono-LP-Rel is rational, Mono-LP-Rel has a rational optimal solution $(y^{*}_{1},\ldots,y^{*}_{k})$.
+To prove the upper bound of the integrality gap, we show that a $k/2$-approximate integer solution to $\mathrm{(LP_{mono})}$ can be constructed from an optimal solution to $\mathrm{(LP_{mono})}$.
+Since the coefficient matrix of the constraints in $\mathrm{(LP_{mono})}$ is rational, $\mathrm{(LP_{mono})}$ has a rational optimal solution $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$.
 Let $\mathcal{L}_{i}:=\{S\subseteq N\mid y^{*}_{i}(S)>0\}$ and $d_{i}:=|\mathcal{L}_{i}|$ for each $i\in[k]$.
@@
 Let $C_{i}^{1}\supsetneq C_{i}^{2}\supsetneq\cdots\supsetneq C_{i}^{d_{i}}$ be the members of $\mathcal{L}_{i}$ for each $i\in[k]$.
-Let $M$ be the product of k(k1)k(k-1) and the denominators of $y^{*}_{i}(C_{i}^{j})$ for all $i,j$.
-Then, $M\cdot y^{*}_{i}$ is an integer vector for every $i\in[k]$.
-For each $i\in[k]$ and $j\in[M]$, define $U_{i}^{j}\subseteq N$ as follows:
-$U_{i}^{j}:=\begin{cases}C_{i}^{\ell}&\mathrm{if\ }1+\displaystyle\sum_{p=1}^{\ell-1}M\cdot y_{i}^{*}(C_{i}^{p})\leq j\leq\displaystyle\sum_{p=1}^{\ell}M\cdot y_{i}^{*}(C_{i}^{p})\ \mathrm{for\ some\ }\ell\in[d_{i}],\\ \emptyset&\mathrm{if\ }1+\displaystyle\sum_{p=1}^{d_{i}}M\cdot y_{i}^{*}(C_{i}^{p})\leq j\leq M.\end{cases}$
-Intuitively, for each $i\in[k]$, we consider the multiset that contains $M\cdot y^{*}_{i}(C_{i}^{\ell})$ copies of $C_{i}^{\ell}$ for all $\ell\in[d_{i}]$, and $U_{i}^{j}$ is the $j$-th largest set in the multiset for each $j\in[M]$.
+For convenience of discussion, we will convert $y^{*}_{i}$ into an integer vector $z_{i}$ by multiplying it by some constant factor, and make $z_{i}(C_{i}^{\ell})$ copies of $C_{i}^{\ell}$ for each $\ell$.
+To this end, we define $M$ as the product of k(k1)k(k-1) and the denominators of $y^{*}_{i}(C_{i}^{j})$ for all $i,j$.
+Then, $z_{i}:=M\cdot y^{*}_{i}$ is an integer vector for every $i\in[k]$.
+For each $i\in[k]$, we define $U_{i}^{1},U_{i}^{2},\ldots,U_{i}^{M}\subseteq N$ by setting $U_{i}^{j}:=C_{i}^{1}$ for the first $z_{i}(C_{i}^{1})$ indices $j=1,2,\ldots,z_{i}(C_{i}^{1})$, and then setting $U_{i}^{j}:=C_{i}^{2}$ for the next $z_{i}(C_{i}^{2})$ indices $j=z_{i}(C_{i}^{1})+1,z_{i}(C_{i}^{1})+2,\ldots,z_{i}(C_{i}^{1})+z_{i}(C_{i}^{2})$, and continue in this way up to $C_{i}^{d_{i}}$.
+For the remaining indices $j$, setting $U_{i}^{j}:=\emptyset$.
+In other words, for each $\ell\in[d_{i}]$, we set $U_{i}^{j}:=C_{i}^{\ell}$ for all $j$ in the following block of indices:
+$1+\displaystyle\sum_{p=1}^{\ell-1}z_{i}(C_{i}^{p})\leq j\leq\displaystyle\sum_{p=1}^{\ell}z_{i}(C_{i}^{p}).$
 By the definition of $U_{i}^{j}$, we have the following lemma.
 Lemma 2.1.
-For each $i\in[k]$ and $\ell=1,\ldots,d_{i}$, the number of $j\in[M]$ with $U_{i}^{j}=C_{i}^{\ell}$ is $M\cdot y_{i}^{*}(C_{i}^{\ell})$.
-We now aim to construct an integer solution to Mono-LP-Rel (i.e., a partition of $N$) from $U_{i}^{j}$s that achieves $k/2$-approximation. To this end, we prepare the following two lemmas.
+For each $i\in[k]$ and $\ell\in[d_{i}]$, the number of $j\in[M]$ such that $U_{i}^{j}=C_{i}^{\ell}$ is $z_{i}(C_{i}^{\ell})$.
+To show that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$, we will construct a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ that corresponds to a $k/2$-approximate solution to $\mathrm{(LP_{mono})}$.
+Such a partition can be constructed by taking the top $2/k$ fraction of the largest sets among $U_{i}^{j}$ for each $i\in[k]$, and partitioning them into groups of $k$ sets whose union is $N$.
+For the construction, we prepare the following two lemmas.
 Lemma 2.2.
 We have
-$\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j})=\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S).$
+$\sum_{i\in[k],j\in[M]}f_{i}(U_{i}^{j})=\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S).$
 Proof.
-By Lemma 2.1, we have
-$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{S\in\mathcal{L}_{i},i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{1\leq\ell\leq d_{i},i\in[k]}y^{*}_{i}(C_{i}^{\ell})f_{i}(C_{i}^{\ell})$
-$=\sum_{1\leq\ell\leq d_{i},i\in[k]}\dfrac{1}{M}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|\cdot f_{i}(C_{i}^{\ell})$
-$=\sum_{\begin{subarray}{c}i\in[k],j\in[M]\\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }1\leq\ell\leq d_{i}\end{subarray}}\dfrac{1}{M}f_{i}(U_{i}^{j})=\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j}).$
-The last equality follows since if $U_{i}^{j}\neq C_{i}^{\ell}$ for all $\ell=1,\ldots,d_{i}$, then we have $U_{i}^{j}=\emptyset$ by the definition. ∎
+Since $z_{i}(C_{i}^{\ell})$ is the number of $j\in[M]$ with $U_{i}^{j}=C_{i}^{\ell}$ by Lemma 2.1, we have
+$\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)=\sum_{S\in\mathcal{L}_{i},i\in[k]}z_{i}(S)f_{i}(S)=\sum_{\ell\in[d_{i}],i\in[k]}z_{i}(C_{i}^{\ell})f_{i}(C_{i}^{\ell})$
+$=\sum_{\ell\in[d_{i}],i\in[k]}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|\cdot f_{i}(C_{i}^{\ell})$
+$=\sum_{\begin{subarray}{c}i\in[k],j\in[M]\\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\end{subarray}}f_{i}(U_{i}^{j})=\sum_{i\in[k],j\in[M]}f_{i}(U_{i}^{j}).$
+The last equality follows since if $U_{i}^{j}\neq C_{i}^{\ell}$ for all $\ell=1,2,\ldots,d_{i}$, then we have $U_{i}^{j}=\emptyset$ by the definition. ∎
 Lemma 2.3.
-Let $a_{1},\ldots,a_{k}\in[M]$ be positive integers such that $\sum_{i=1}^{k}a_{i}\leq M+k-1$. Then, $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$.
+Let $a_{1},a_{2},\ldots,a_{k}\in[M]$ be positive integers such that $\sum_{i=1}^{k}a_{i}\leq M+k-1$. Then, $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$.
 Proof.
@@
 Then, there exists $e\in N$ such that $e\notin U_{i}^{a_{i}}$ for every $i\in[k]$.
-Combined with $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ for each $i\in[k]$, this implies $e\notin U_{i}^{j}$ for every $i\in[k]$ and integer $j$ with $a_{i}\leq j\leq M$.
-Hence, by Lemma 2.1 we have
-$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(S)=\sum_{\begin{subarray}{c}S\in\mathcal{L}_{i},e\in S\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(S)=\sum_{\begin{subarray}{c}1\leq\ell\leq d_{i},e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(C_{i}^{\ell})$
-$=\sum_{\begin{subarray}{c}1\leq\ell\leq d_{i},e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|$
-$=\sum_{i\in[k]}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}],\ e\in U_{i}^{j}\right\}\right|$
-$\leq\sum_{i\in[k]}\left|\left\{j\in[a_{i}-1]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\right\}\right|$
+Combined with $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ for each $i\in[k]$, this implies that $e\notin U_{i}^{j}$ for every $i\in[k]$ and integer $j$ with $a_{i}\leq j\leq M$.
+Hence, since $z_{i}(C_{i}^{\ell})$ is the number of $j\in[M]$ with $U_{i}^{j}=C_{i}^{\ell}$ by Lemma 2.1, we have
+$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=\sum_{\begin{subarray}{c}S\in\mathcal{L}_{i},e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=\sum_{\begin{subarray}{c}\ell\in[d_{i}],e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}z_{i}(C_{i}^{\ell})$
+$=\sum_{\begin{subarray}{c}\ell\in[d_{i}],e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|$
+$=\sum_{i\in[k]}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\ \mathrm{with\ }e\in C_{i}^{\ell}\right\}\right|$
+$=\sum_{i\in[k]}\left|\left\{j\in[a_{i}-1]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\ \mathrm{with\ }e\in C_{i}^{\ell}\right\}\right|$
 $\leq\sum_{i\in[k]}(a_{i}-1)\leq M-1.$ (2.1)
-Since $y^{*}_{1},\ldots,y^{*}_{k}$ is a feasible solution to Mono-LP-Rel, we have
+Since $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$ is a feasible solution to $\mathrm{(LP_{mono})}$, we have
 $\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y_{i}^{*}(S)=1.$
 This implies
-$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}M\cdot y_{i}^{*}(S)=M,$
+$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=M,$
 which contradicts the inequality (2). ∎
 Since $M$ is divisible by k(k1)k(k-1), we can denote $M=k(k-1)\cdot m$ for some integer $m$.
-For $k$ positive integers $a_{1},\ldots,a_{k}\in[M]$, we say that an ordered tuple $(a_{1},\ldots,a_{k})$ covers $N$ if $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$.
-We show that if there exist distinct $M\cdot 2/k=2m(k-1)$ ordered tuples covering $N$, then one of them corresponds to an integer solution to Mono-LP-Rel that achieves $k/2$-approximation.
+For $k$ positive integers $a_{1},a_{2},\ldots,a_{k}\in[M]$, we say that an ordered tuple $(a_{1},a_{2},\ldots,a_{k})$ covers $N$ if $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$.
+Note that if an ordered tuple $(a_{1},a_{2},\ldots,a_{k})$ covers $N$, then there exist $S_{i}\subseteq U_{i}^{a_{i}}$ for $i=1,2,\ldots,k$ such that $(S_{1},S_{2},\ldots,S_{k})$ is a partition of $N$ and $\sum_{i=1}^{k}f_{i}(S_{i})\leq\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}})$.
+We first show that if there exist 2m(k1)2m(k-1) pairwise disjoint ordered tuples covering $N$, then one of them induces a partition $(S_{1},S_{2},\ldots,S_{k})$ of $N$ that corresponds to a $k/2$-approximate solution to $\mathrm{(LP_{mono})}$.
+We then construct 2m(k1)2m(k-1) such pairwise disjoint ordered tuples covering $N$, which completes the proof of Theorem 1.1.
+The following lemma is the first part of the proof:
 Lemma 2.4.
@@
   • (1)
-    $a_{i}^{j}\in[M]$ for each $i\in[k]$ and $j\in[2m(k-1)]$,
+    $a_{i}^{j}\in[M]$ for every $i\in[k]$ and $j\in[2m(k-1)]$,
   • (2)
-    $a_{i}^{j}\neq a_{i}^{\ell}$ for each $i\in[k]$ and $j,\ell\in[2m(k-1)]$ with $j\neq\ell$,
+    $a_{i}^{j}\neq a_{i}^{\ell}$ for every $i\in[k]$ and $j,\ell\in[2m(k-1)]$ with $j\neq\ell$,
   • (3)
-    each ordered tuple covers $N$.
-Then, the integrality gap of Mono-LP-Rel is at most $k/2$.
+    every ordered tuple covers $N$.
+Then, the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$.
 Proof.
 By Lemma 2.2, we have
-$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j})\geq\sum_{i\in[k],j\in[2m(k-1)]}\dfrac{1}{M}f_{i}(U_{i}^{a_{i}^{j}})$
-$\geq\dfrac{2m(k-1)}{M}\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}$
-$=\dfrac{2}{k}\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}.$ (2.2)
+$\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)=\sum_{i\in[k],j\in[M]}f_{i}(U_{i}^{j})\geq\sum_{i\in[k],j\in[2m(k-1)]}f_{i}(U_{i}^{a_{i}^{j}})$
+$\geq 2m(k-1)\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}.$ (2.2)
 The first inequality follows since by the monotonicity of $f_{i}$, we have $f_{i}(U_{i}^{j})\geq f_{i}(\emptyset)=0$ for every $i\in[k]$ and $j\in[M]$. Let
@@
 Then, by (2) we have
-$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\dfrac{2}{k}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$
-Hence, we have
+$\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)\geq 2m(k-1)\cdot\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$
+Since $z_{i}=M\cdot y^{*}_{i}=mk(k-1)\cdot y^{*}_{i}$ for each $i\in[k]$, this implies
 $\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$ (2.3)
-Since $\bigcup_{i=1}^{k}U_{i}^{a_{i}^{j^{*}}}=N$, there exists a partition $U_{1},\ldots,U_{k}$ of $N$ such that $U_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$.
+Since $\bigcup_{i=1}^{k}U_{i}^{a_{i}^{j^{*}}}=N$, there exists a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ such that $S_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$.
 Then, by (2.3) and the monotonicity of $f_{i}$, we have
-$\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}})\geq\sum_{i\in[k]}f_{i}(U_{i}).$
-This implies that the integrality gap of Mono-LP-Rel is at most $k/2$. ∎
-We are now ready to prove Theorem 1.1.
-Theorem 1.1.
+$\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}})\geq\sum_{i\in[k]}f_{i}(S_{i}).$
+This implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$. ∎
+We now prove Theorem 1.1 by constructing 2m(k1)2m(k-1) pairwise disjoint ordered tuples covering $N$, using only the smallest $2/k$ fraction of the indices $j=1,2,\ldots,M$ of $U_{i}^{j}$ for each $i\in[k]$.
+Proof of Theorem 1.1.
 By Lemma 2.4, it suffices to show that there exist 2m(k1)2m(k-1) ordered tuples satisfying the conditions in Lemma 2.4.
-For each positive integer $a\in\mathbb{Z}_{>0}$, define $\mathrm{mod}(a)$ as follows:
-[…]
+For each positive integer $a\in\mathbb{Z}_{>0}$, define $\mathrm{mod}(a)$ to be the unique integer $r\in[2m(k-1)]$ such that $a\equiv r\pmod{2m(k-1)}$.
 For each $i\in[k]$ and $j\in[2m(k-1)]$, define $a_{i}^{j}$ as follows:
@@
 We show that $(a_{i}^{j})_{i\in[k],j\in[2m(k-1)]}$ satisfies conditions (1)–(3) in Lemma 2.4.
-For condition (1), we show that $1\leq a_{i}^{j}\leq M$ for each $i\in[k]$ and $j\in[2m(k-1)]$.
+For condition (1), we show that $a_{i}^{j}\in[M]$ for each $i\in[k]$ and $j\in[2m(k-1)]$.
 If $1\leq i\leq k-1$, then we have
@@
 This completes the proof. ∎
-The proof of Theorem 1.1 provides the construction of a $k/2$-approximate integer solution to Mono-LP-Rel.
-We prove Theorem 1.2 by showing that such an integer solution can be constructed in polynomial time.
-Theorem 1.2.
-We can obtain a rational optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to Mono-LP-Rel whose supports $\mathcal{L}_{1},\ldots,\mathcal{L}_{k}$ are chains in polynomial time via the ellipsoid method [5] (see Appendix A for details).
+The proofs of Lemma 2.4 and Theorem 1.1 provide the construction of a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ that corresponds to a $k/2$-approximate solution to $\mathrm{(LP_{mono})}$, and hence it is also a $k/2$-approximate solution to MMSCA.
+The construction works as follows.
+We first compute a rational optimal solution $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$ to $\mathrm{(LP_{mono})}$ whose supports are chains, and define $U_{i}^{j}$ for each $i\in[k]$ and $j\in[M]$.
+We then define 2m(k1)2m(k-1) ordered tuples $(a_{1}^{j},a_{2}^{j},\ldots,a_{k}^{j})$ covering $N$ as in the proof of Theorem 1.1, and compute the index $j^{*}$ minimizing $\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}^{j}})$ in the proof of Lemma 2.4.
+Finally, we obtain sets $U_{1}^{a_{1}^{j^{*}}},U_{2}^{a_{2}^{j^{*}}},\ldots,U_{k}^{a_{k}^{j^{*}}}$, which induces a desired partition $S_{1},S_{2},\ldots,S_{k}$ of $N$.
+We prove Theorem 1.2 by showing that the above construction can be done in polynomial time.
+Proof of Theorem 1.2.
+We can obtain a rational optimal solution $(y_{1}^{*},y_{2}^{*},\ldots,y_{k}^{*})$ to $\mathrm{(LP_{mono})}$ whose supports $\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{k}$ are chains in polynomial time via the ellipsoid method [5] (see Appendix A for details).
 Then, we define $U_{i}^{j}$ for each $i\in[k]$ and $j\in[M]$, and $a_{i}^{j}$ for each $i\in[k]$ and $j\in[2m(k-1)]$ as in the proof of Theorem 1.1.
 If we can compute
-$j^{*}:=\underset{j\in[2m(k-1)]}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
-then any partition $U_{1},\ldots,U_{k}$ of $N$ such that $U_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$ satisfies
-$\sum_{i\in[k]}f_{i}(U_{i})\leq\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\dfrac{k}{2}\cdot\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\},$
+$j^{*}\in\underset{j\in[2m(k-1)]}{\operatorname{argmin}}\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}^{j}}),$
+then any partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ such that $S_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$ satisfies
+$\sum_{i=1}^{k}f_{i}(S_{i})\leq\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}^{j^{*}}})\leq\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$
+$\leq\dfrac{k}{2}\cdot\min\left\{\sum_{i=1}^{k}f_{i}(T_{i})\ \middle|\ T_{1},T_{2},\ldots,T_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\},$
 as desired.
@@
 Then, for q=2,3,,q=2,3,..., we define $j_{q}$ as the minimum integer such that $j_{q}>j_{q-1}$ and there exists $i\in[k]$ with $U_{i}^{a_{i}^{j_{q}}}\neq U_{i}^{a_{i}^{j_{q-1}}}$.
-Recall that $n:=|N|$.
+Recall that $n=|N|$.
 Since for each $i\in[k]$, we have $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ and there are at most $|\mathcal{L}_{i}|\leq n$ distinct sets among $\{U_{i}^{j}\}_{j=1}^{M}$, we only need to define $j_{q}$ for $q=1,2,\ldots,Q$ with $Q\leq kn$.
 Since we have
-$j^{*}=\underset{j\in\{j_{1},j_{2},\ldots,j_{Q}\}}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
+$j^{*}\in\underset{j\in\{j_{1},j_{2},\ldots,j_{Q}\}}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
 we can find $j^{*}$ by computing the values of $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$.
-By the definition of $U_{i}^{j}$ and $a_{i}^{j}$, we can compute $j_{q}$ in polynomial time for each fixed $q$.
+By the definition of $U_{i}^{j}$ and $a_{i}^{j}$, we can compute $j_{q}$ in polynomial time for each $q\in[Q]$.
 Hence, since $Q\leq kn$, we can compute $j_{q}$ and $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$ in polynomial time.
@@
 3 Lower bound on the integrality gap
-In this section, we show that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$.
-To this end, we explicitly construct monotone submodular functions $f_{1},\ldots,f_{k}$ such that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
+In this section, we show that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$.
+To this end, we explicitly construct monotone submodular functions $f_{1},f_{2},\ldots,f_{k}$ such that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
 Suppose that $\log n\geq k$. Let $p$ be the unique positive integer such that $pk\leq\lfloor\log n\rfloor\leq(p+1)k-1$. We define
@@
 The last inequality follows from the monotonicity of $f_{i}$. Therefore, $f_{i}$ is submodular. ∎
-We now show that for the monotone submodular functions $f_{1},\ldots,f_{k}$ defined above, any optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to Mono-LP-Rel satisfies
+We now show that for the monotone submodular functions $f_{1},f_{2},\ldots,f_{k}$ defined above, any optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to $\mathrm{(LP_{mono})}$ satisfies
 $\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$
-$\leq\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$ (3.1)
-The inequality (3) implies that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$, which completes the proof of Theorem 1.3.
+$\leq\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$ (3.1)
+The inequality (3) implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$, which completes the proof of Theorem 1.3.
 To prove (3), we first give an upper bound on the value $\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$.
-To this end, we construct a feasible solution to Mono-LP-Rel.
+To this end, we construct a feasible solution to $\mathrm{(LP_{mono})}$.
 For each $i\in[k]$ and $j\in[pk-k+1]$, let $N_{i}^{j}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}\geq j\}$.
@@
 Let $N^{c}:=N\setminus\{e_{v}\in N\mid v\in\mathcal{V}\}$.
-For each $i\in[k]$, define $z_{i}\in\mathbb{R}_{+}^{2^{N}}$ as follows: for all $S\subseteq N$,
-$z_{i}(S):=\begin{cases}\dfrac{1}{pk-k+1}&\mathrm{if\ }i=1\ \mathrm{and\ }\vskip 3.0ptS=N_{1}^{j}\cup N^{c}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ \vskip 3.0pt\dfrac{1}{pk-k+1}&\mathrm{if\ }2\leq i\leq k\ \mathrm{and\ }S=N_{i}^{j}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ 0&\mathrm{otherwise}.\end{cases}$
+For each $i\in[k]$, define $y^{\prime}_{i}\in\mathbb{R}_{+}^{2^{N}}$ as follows: for all $S\subseteq N$,
+$y^{\prime}_{i}(S):=\begin{cases}\dfrac{1}{pk-k+1}&\mathrm{if\ }i=1\ \mathrm{and\ }\vskip 3.0ptS=N_{1}^{j}\cup N^{c}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ \vskip 3.0pt\dfrac{1}{pk-k+1}&\mathrm{if\ }2\leq i\leq k\ \mathrm{and\ }S=N_{i}^{j}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ 0&\mathrm{otherwise}.\end{cases}$
 Lemma 3.2.
-$(z_{1},\ldots,z_{k})$ is a feasible solution to Mono-LP-Rel.
+$(y^{\prime}_{1},y^{\prime}_{2},\ldots,y^{\prime}_{k})$ is a feasible solution to $\mathrm{(LP_{mono})}$.
 Proof.
-To show the feasibility of $(z_{1},\ldots,z_{k})$, we show that
-$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=1$ (3.2)
+To show the feasibility of $(y^{\prime}_{1},y^{\prime}_{2},\ldots,y^{\prime}_{k})$, we show that
+$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y^{\prime}_{i}(S)=1$ (3.2)
 for every $e\in N$. For any $e\in N^{c}$, we have
-$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=\sum_{j\in[pk-k+1]}z_{1}(N_{1}^{j}\cup N^{c})=\sum_{j\in[pk-k+1]}\dfrac{1}{pk-k+1}=1.$
+$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y^{\prime}_{i}(S)=\sum_{j\in[pk-k+1]}y^{\prime}_{1}(N_{1}^{j}\cup N^{c})=\sum_{j\in[pk-k+1]}\dfrac{1}{pk-k+1}=1.$
 Take any $e\in N\setminus N^{c}$. Since $e=e_{v}$ for some $v\in\mathcal{V}$, we have
-$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)$ $=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}z_{1}(N_{1}^{j}\cup N^{c})+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}z_{i}(N_{i}^{j})$
+$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y^{\prime}_{i}(S)$ $=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}y^{\prime}_{1}(N_{1}^{j}\cup N^{c})+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}y^{\prime}_{i}(N_{i}^{j})$
 $=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}\frac{1}{pk-k+1}+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}\frac{1}{pk-k+1}$
@@
 $=\frac{pk-k+1}{pk-k+1}=1.$
-The second last equality follows since $\sum_{i=1}^{k}v_{i}=pk-k+1$. Hence, (3.2) holds for every $e\in N$, which implies that $(z_{1},\ldots,z_{k})$ is a feasible solution to Mono-LP-Rel. ∎
-By Lemma 3.2 and the optimality of $y^{*}_{1},\ldots,y^{*}_{k}$, we have
-$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)$
+The second last equality follows since $\sum_{i=1}^{k}v_{i}=pk-k+1$.
+Hence, (3.2) holds for every $e\in N$, which implies that $(y^{\prime}_{1},y^{\prime}_{2},\ldots,y^{\prime}_{k})$ is a feasible solution to $\mathrm{(LP_{mono})}$.
+∎
+By Lemma 3.2 and the optimality of $y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k}$, we have
+$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\sum_{S\subseteq N,i\in[k]}y^{\prime}_{i}(S)f_{i}(S)$
 $=\sum_{j\in[pk-k+1]}\frac{1}{pk-k+1}f_{1}(N_{1}^{j}\cup N^{c})+\sum_{2\leq i\leq k,j\in[pk-k+1]}\frac{1}{pk-k+1}f_{i}(N_{i}^{j})$
@@
 Lemma 3.3.
-Let $X_{1},\ldots,X_{k}$ be any partition of $N$. Then, we have
-$\sum_{i=1}^{k}f_{i}(X_{i})\geq pk+k.$
+Let $S_{1},S_{2},\ldots,S_{k}$ be any partition of $N$. Then, we have
+$\sum_{i=1}^{k}f_{i}(S_{i})\geq pk+k.$
 Proof.
-If $X_{j}\cap Z_{j}\neq\emptyset$ for some $j\in[k]$, then we have
-$\sum_{i=1}^{k}f_{i}(X_{i})\geq f_{j}(X_{j})=2pk+1\geq pk+k.$
-Hence, it suffices to consider the case when $X_{i}\cap Z_{i}=\emptyset$ for every $i\in[k]$. For each $i\in[k]$, define $x^{i}\in\mathcal{V}$ as follows:
+If $S_{j}\cap Z_{j}\neq\emptyset$ for some $j\in[k]$, then we have
+$\sum_{i=1}^{k}f_{i}(S_{i})\geq f_{j}(S_{j})=2pk+1\geq pk+k.$
+Hence, it suffices to consider the case when $S_{i}\cap Z_{i}=\emptyset$ for every $i\in[k]$. For each $i\in[k]$, define $x^{i}\in\mathcal{V}$ as follows:
 $x^{i}_{j}:=\begin{cases}pk-k+1&\mathrm{if\ }j=i,\\ 0&\mathrm{if\ }j\in[k]\mathrm{\ and\ }j\neq i.\end{cases}$
-Then, since $X_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have $e_{x^{i}}\in X_{i}$ for every $i\in[k]$.
-Hence, we have $X_{i}\cap N_{i}\neq\emptyset$ for every $i\in[k]$.
-For each $i\in[k]$, we define $b_{i}:=\min\{v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}$.
+Then, since $S_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have $e_{x^{i}}\in S_{i}$ for every $i\in[k]$.
+Hence, we have $S_{i}\cap N_{i}\neq\emptyset$ for every $i\in[k]$.
+For each $i\in[k]$, we define $b_{i}:=\min\{v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}$.
 We now show that $\sum_{i=1}^{k}b_{i}\leq pk$.
@@
 Then, since $\sum_{i=1}^{k}(b_{i}-1)\geq pk-k+1$, there exists $v\in\mathcal{V}$ such that $v_{i}\leq b_{i}-1$ for every $i\in[k]$.
-Choose $i\in[k]$ such that $e_{v}\in X_{i}$.
-Since $X_{i}\cap Z_{i}=\emptyset$ by the assumption, we have $e_{v}\in X_{i}\cap N_{i}$, which contradicts the definition of $b_{i}$.
+Choose $i\in[k]$ such that $e_{v}\in S_{i}$.
+Since $S_{i}\cap Z_{i}=\emptyset$ by the assumption, we have $e_{v}\in S_{i}\cap N_{i}$, which contradicts the definition of $b_{i}$.
 Hence, we have $\sum_{i=1}^{k}b_{i}\leq pk$.
-Therefore, since $X_{i}\cap N_{i}\neq\emptyset$ and $X_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have
-$\sum_{i=1}^{k}f_{i}(X_{i})=\sum_{i=1}^{k}\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}\}$
-$\geq\sum_{i=1}^{k}\max\{2p+1-v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}=2pk+k-\sum_{i=1}^{k}\min\{v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}$
+Therefore, since $S_{i}\cap N_{i}\neq\emptyset$ and $S_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have
+$\sum_{i=1}^{k}f_{i}(S_{i})=\sum_{i=1}^{k}\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}\}$
+$\geq\sum_{i=1}^{k}\max\{2p+1-v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}=2pk+k-\sum_{i=1}^{k}\min\{v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}$
 $=2pk+k-\sum_{i=1}^{k}b_{i}\geq 2pk+k-pk=pk+k.$
 By Lemma 3.3, we have
-$\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}$
+$\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}$
 $\geq pk+k=\frac{2k(p+1)^{2}(pk-k+1)}{2(p+1)(pk-k+1)}$
@@
 $\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\frac{pk(2p+1)}{pk-k+1}$
-$\leq\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
+$\leq\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
 Thus, the inequality (3) holds, which completes the proof of Theorem 1.3.
@@
 Appendix A Convex program based on the Lovász extension
-In this section, we show that LP-Rel is equivalent to the convex program relaxation of MSCA introduced by Chekuri and Ene [5].
+In this section, we show that (LP) is equivalent to the convex program relaxation of MSCA introduced by Chekuri and Ene [5].
 Let $N:=\{1,\ldots,n\}$.
 Let $f_{i}:2^{N}\to\mathbb{R}$ be a submodular function with $f_{i}(\emptyset)=0$ for each $i\in[k]$.
-Recall that LP-Rel is defined as follows:
-$(\textsc{LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
+Recall that (LP) is defined as follows:
+$\textsc{(LP)}\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
 $\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
@@
 $\ \ \ \sum_{i=1}^{k}x^{i}_{j}=1\ \mathrm{for\ every\ }j\in[n].$
-We now show that any feasible solution $({x^{1}},\ldots,x^{k})$ to LE-Rel can be translated into a feasible solution $(y_{1},\ldots,y_{k})$ to LP-Rel that has the same objective value and whose supports are chains.
-Take any feasible solution $({x^{1}},\ldots,x^{k})$ to LE-Rel.
+We now show that any feasible solution $({x^{1}},\ldots,x^{k})$ to (LE-Rel) can be translated into a feasible solution $(y_{1},y_{2},\ldots,y_{k})$ to (LP) that has the same objective value and whose supports are chains.
+Take any feasible solution $({x^{1}},\ldots,x^{k})$ to (LE-Rel).
 We will describe a translation from $x^{i}$ to $y_{i}$ for each $i\in[k]$.
@@
 $y_{i}(S):=\begin{cases}x_{j}^{i}-x_{j+1}^{i}&\mathrm{if\ }S=S_{j}\ \mathrm{and\ }x_{j}^{i}-x_{j+1}^{i}>0\ \mathrm{for\ some\ }j\in[n],\\ 0&\mathrm{otherwise}.\end{cases}$
-Then, $(y_{1},\ldots,y_{k})$ is a feasible solution to LP-Rel because for each $i\in[k]$ we have
+Then, $(y_{1},y_{2},\ldots,y_{k})$ is a feasible solution to (LP) because for each $i\in[k]$ we have
 $\sum_{S\subseteq N}y_{i}(S)\chi_{S}=\sum_{j\in[n]}(x^{i}_{j}-x^{i}_{j+1})\chi_{S_{j}}=x^{i}$
@@
 $\sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)=\sum_{i=1}^{k}\hat{f_{i}}(x^{i}),$
-which implies that $(y_{1},\ldots,y_{k})$ has the same objective value as $(x^{1},\ldots,x^{k})$.
-Similarly, we can translate any feasible solution to LP-Rel whose supports are chains into a feasible solution to LE-Rel that has the same objective value.
-Since LE-Rel can be solved in polynomial time via the ellipsoid method [5], an optimal solution to LP-Rel whose supports are chains can be obtained in polynomial time.
+which implies that $(y_{1},y_{2},\ldots,y_{k})$ has the same objective value as $(x^{1},\ldots,x^{k})$.
+Similarly, we can translate any feasible solution to (LP) whose supports are chains into a feasible solution to (LE-Rel) that has the same objective value.
+Since (LE-Rel) can be solved in polynomial time via the ellipsoid method [5], an optimal solution to (LP) whose supports are chains can be obtained in polynomial time.
 Chekuri and Ene [6] provided another equivalent linear programming formulation of MSCA called LE-Rel-Primal.
-We note that LP-Rel is equivalent to LE-Rel-Primal.
+We note that (LP) is equivalent to LE-Rel-Primal.
 Appendix B $k$-polymatroid intersection
@@
 The linear program $\mathrm{LP}_{k\text{-poly}}$ corresponds to the weighted polymatroid intersection when $k=2$.
-When $w=\mathbf{1}$, the linear program $\mathrm{LP}_{k\text{-poly}}$ is the dual linear program of LP-Rel.
-For the case when $f_{1},\ldots,f_{k}$ are matroid rank functions (i.e., the case of weighted $k$-matroid intersection), the integrality gap of $\mathrm{LP}_{k\text{-poly}}$ is $2$ when $k=3$ [13], and it equals k1k-1 for arbitrary $k$ in the unweighted case $w=\mathbf{1}$ [12].
+When $w=\mathbf{1}$, the linear program $\mathrm{LP}_{k\text{-poly}}$ is the dual linear program of (LP).
+For the case when $f_{1},f_{2},\ldots,f_{k}$ are matroid rank functions (i.e., the case of weighted $k$-matroid intersection), the integrality gap of $\mathrm{LP}_{k\text{-poly}}$ is $2$ when $k=3$ [13], and it equals k1k-1 for arbitrary $k$ in the unweighted case $w=\mathbf{1}$ [12].